# Model averaging with mixed criteria for estimating high quantiles of extreme values: Application to heavy rainfall

### v1

@@ 2.3 Estimation of high quantiles and building a surrogate model
which is a weighted average of the return level predictions from individual submodels. We aim to achieve this with lower uncertainty than the MLE, LME, and other established methods.
The limitation of this MA approach is that it may not be useful for system description and understanding (Fletcher 2018). We consider constructing a weighted averaged (or ensemble) model from the $K$ GEV candidate models to address this limitation. However, the ensemble model may not necessarily follow a GEVD, even though each model $M_{k}$ does. Thus, simply taking a weighted sum of $K$ GEV candidate models is not an ideal solution. Instead, we propose building a surrogate model that still follows a GEVD and best approximates the return level estimates generated by the MA method.
@@
where $\{r_{q_{i}}\}_{i=1,\dots,I}$ are the return levels, corresponding to probability $q_{i}$, computed from a $GEVD(\mu,\sigma,\xi)$, and $\hat{r}_{q_{i}}$ are the corresponding return level estimates obtained using the MA method. To ensure a broad coverage of important quantiles, we selected 16 probabilities ranging from $0.5$ to $0.999$. Although the surrogate model is only an approximation and may not be theoretically exact, it is computationally convenient and useful for further data analysis. This convenience includes additional predictions, model comparison, and model diagnostics such as quantile-per-quantile and return level plots. Of course, other models could also be used as surrogates instead of reusing the GEVD.
@@ 3.1 Conditional expectation
We provide a theoretical justification for why the proposed MA method is expected to perform well. If we assume that $r(\hat{\mu},\hat{\sigma},\hat{\xi})$ is an unbiased estimator of the return level $r(\mu,\sigma,\xi)$, then we have:
@@
If $E\{r(\hat{\mu},\hat{\sigma},\hat{\xi})|\hat{\xi}=\xi\}$ serves as an estimator of $r(\mu,\sigma,\xi)$, then it is unbiased and has a smaller variance than $r(\hat{\mu},\hat{\sigma},\hat{\xi})$, as follows from the conditional variance identity (Casella & Berger 2001),
$Var\{r(\hat{\mu},\hat{\sigma},\hat{\xi})\}$ $\ =\ Var\left[E\{r(\hat{\mu},\hat{\sigma},\hat{\xi})\>|\>\hat{\xi}=\xi\}\right]\>+\>E\left[Var\{r(\hat{\mu},\hat{\sigma},\hat{\xi})\>|\>\hat{\xi}=\xi\}\right]$ (11)
@@
If $\hat{\xi}$ is a sufficient statistic for $r(\mu,\sigma,\xi)$, then $E\{r(\hat{\mu},\hat{\sigma},\hat{\xi})\,|\,\hat{\xi}=\xi\}$ is a uniformly better-unbiased estimator, as stated by the Rao-Blackwell theorem (Casella & Berger 2001). However, $E\{r(\hat{\mu},\hat{\sigma},\hat{\xi})\,|\,\hat{\xi}=\xi\}$ is not an estimator, as it depends on the unknown parameter $\xi$. To address this, we approximate $E\{r(\hat{\mu},\hat{\sigma},\hat{\xi})\,|\,\hat{\xi}=\xi\}$ using a weighted average of available estimators computed for values of $\xi$s near $\hat{\xi}$, leveraging relatively simple estimation methods.
Since estimating $\xi$ is more challenging than estimating the other parameters of the GEVD and because $\hat{\xi}$ has the most substantial influence on the estimation of $r(\mu,\sigma,\xi)$—particularly when $\xi<-0.2$—conditioning $\xi$ as a fixed value near $\hat{\xi}$ can substantially simplify the estimation process. The generally lower variability of a narrow submodel compared to a broader model is well demonstrated in Claeskens & Hjort (2008) under the local misspecification framework. Additionally, fitting submodels is computationally more efficient than fitting a full model.
Assuming that $(\hat{\mu},\hat{\sigma},\hat{\xi})$ can always be estimated from the data, a heuristic derivation of $\hat{r}_{\text{MA}}$ proceeds as follows:
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\begin{pmatrix}\hat{\mu}_{2}(\xi),\hat{\sigma}_{2}(\xi)\text{ are 2-dimentional }\\ \text{estimators for a given constant}\ \xi\end{pmatrix}$
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\begin{pmatrix}\hat{\mu}_{2}(\xi),\>\hat{\sigma}_{2}(\xi)\text{ are no more random}\\ \text{ for a given constant $\xi$ and the data}\end{pmatrix}$
@@
[…]
where $\xi_{k}$ values are selected based on the sampling distribution of $\xi$. Thus, our MA method can be interpreted as integrating out the $\xi$ parameter.
@@ 3.2 Asymptotic distribution of MA return level
When submodel $M_{k}$ is fitted using the MLE, the MLE of $(\mu_{k},\ \sigma_{k})$ for a given $\xi_{k}$ follows asymptotically a normal distribution with mean $(\mu_{k},\;\sigma_{k})$ and covariance matrix $\Sigma_{k}$, which is the inverse of the Fisher information matrix, provided that $-1<\xi_{k}<1/2$ (Bücher et al. 2021, for example). Consequently, $\hat{r}_{k}$ asymptotically follows a normal distribution with mean $r_{k}$ and variance:
@@
by the delta method, where $\nabla r_{k}$ is a gradient vector of $r_{k}$ with respect to $(\mu_{k},\;\sigma_{k})$. For the GEV distribution, the gradient vector of $r_{k}$ with respect to $(\mu_{k},\;\sigma_{k})$ is given by: $\nabla r_{k}=[1,\;\xi_{k}^{-1}(1-y_{p}^{\xi_{k}})]^{T}$, for given $\xi_{k}$, where $\,y_{p}=-\ln(1-p)$.
@@
[…]
@@
[…]
where $\Lambda_{k}\,\sim AN(0,\>v_{k})$.
[…]
@@
[…]
@@
The Slusky theorem and continuity theorem of weak convergence are applied in the above derivation. This establishes the statement of the theorem.∎
@@
where $\>\underline{w}^{T}=(w_{1},\dots,w_{K})$ and $C$ is $Var(\underline{\Lambda})$ which is the asymptotic covariance matrix of ^$\hat{\>}\underline{r}$, where $\underline{\Lambda}^{T}=(\Lambda_{1},\dots,\Lambda_{K})$.
@@ 3.3 Asymptotic variance of MA return level
The diagnal elements of $C$, the asymptotic variances of ^$\hat{\>}\underline{r}$, are obtained using (13). However, the off-diagonal elements of $C$, representing the covariances $\text{Cov}(\hat{r}_{i},\;\hat{r}_{j})$ for $i!=j$, are difficult to compute explicitly. To address this, we considered a variance formula directly from (8):
@@ 3.4 MA variance with random weights
For computing $Var(\Lambda_{MA})$ in real applications with data of size $n$, we may need to replace $\>\underline{w}$ in (18) with its estimate ^$\hat{\>}\underline{w}^{T}=(\hat{w}_{1},\dots,\hat{w}_{K})$. In this case, $\>\hat{\underline{w}}$ is a random vector derived from the sample, meaning that the formula (18) may need to be adjusted to account for the uncertainty in $\>\hat{\underline{w}}$. To calculate $Var(\Lambda_{MA})$ using the random vector $\>\hat{\underline{w}}$, we assume that $\>\hat{\underline{w}}$ and ^$\hat{\>}\underline{r}$ are independent. The expression for the variance is then given by:
@@
$=\ Var_{\>\hat{\underline{w}}}\left[E(\>\hat{\underline{w}}^{T}\hat{\>}\underline{r}\,|\,\>\hat{\underline{w}})\right]\;+\;E_{\>\hat{\underline{w}}}\left[Var(\>\hat{\underline{w}}^{T}\hat{\>}\underline{r}\,|\,\>\hat{\underline{w}})\right]$
@@
$=\ E(\hat{\>}\underline{r})^{T}\,D\,E(\hat{\>}\underline{r})\;+\;tr(D\,C)\;+\;E(\>\hat{\underline{w}})^{T}\,C\,E(\>\hat{\underline{w}}),$
where $D=Var(\>\hat{\underline{w}})$. In the last equation, we used the formula $E(\>\hat{\underline{w}}^{T}\,C\,\>\hat{\underline{w}})=tr(D\,C)+E(\>\hat{\underline{w}})^{T}\,C\,E(\>\hat{\underline{w}})$ (Ravishanker et al. 2022, for example). Here, $C$ is approximated by the asymptotic covariance matrix $(\hat{C})$, computed using (13) and (20), and evaluated at ^$\hat{\>}\underline{r}$. Due to the additional uncertainty introduced by the randomness of the weights, this variance must be greater than or equal to the asymptotic variance in (19) for fixed weights.
We need to estimate $E(\>\hat{\underline{w}})$, $E(\hat{\>}\underline{r})$, and $D$ from the data for practical purposes. However, deriving these theoretically from (7) is challenging. One possible approach is to use bootstrapping to approximate these quantities, although we did not adopt this method in our study. Instead, we assume a distribution for $\>\hat{\underline{w}}$. Given that $0\leq\hat{w}_{k}\leq 1$ for $k=1,\dots,K$ and that weights $\sum_{k=1}^{K}\hat{w}_{k}=1$, it is reasonable to assume that $\>\hat{\underline{w}}$ follows a Dirichlet distribution with parameter vector $\>\underline{w}$, where $0\leq w_{k}\leq 1$ for $k=1,\dots,K$ and $\sum_{k=1}^{K}w_{k}=1$. Then, under this assumption,
@@
for all $i\neq j$. These quantities form the elements of $D$. The covariance matrix $D$ can then be approximated by substituting ^$\hat{\>}\underline{w}$ in place of $\>\underline{w}$, yielding the estimated matrix by $\hat{D}$.
To approximate $E(\hat{\>}\underline{r})$, we used a moving average of order $q$ for ^$\hat{\>}\underline{r}$ instead of directly substituting ^$\hat{\>}\underline{r}$ into $E(\hat{\>}\underline{r})$. With these approximations, we obtain $Var(\hat{r}_{\text{MA}})$ in (21) using the following formulation:
@@
where ^$\widetilde{\hat{\>}\underline{r}(q)}^{T}$ is the order $q$ moving average of ^$\hat{\>}\underline{r}^{T}=(\hat{r}_{1},\dots,\hat{r}_{K})$. We chose $q=3$ in this study.
This formula is computationally much faster than using bootstrap to estimate $E(\>\hat{\underline{w}})$, $E(\hat{\>}\underline{r})$, and $D$. This expression for the MA variance with random weights has not been previously documented in the literature.
@@ 4.1 Computing strategy
Selecting the number ($K$) of candidate models is necessary in the proposed method. One recommended strategy is to start with a relatively large $K$ (e.g., $K=15$) and then remove submodels with small weights, such as those with $w_{k}\leq 0.01$. Additionally, if the weight assigned to the smallest $\xi_{k}$ or the largest $\xi_{k}$ remains above a certain threshold (e.g., 0.2), the method adds a few additional $xi_{k}$ values beyond the lower or upper bounds. The weights are then recalculated using the updated number of submodels, $K^{\prime}$, and the MA return level estimate $\hat{r}_{MA}$ is obtained using the new weight vector ${\>\hat{\underline{w}}^{\prime}}^{T}=(\hat{w}^{\prime}_{i_{1}},\dots,\hat{w}^{\prime}_{i_{K^{\prime}}})$.
Thus, selecting between 90% and 95% confidence intervals for choosing $\xi_{k}$s values may no longer be critically important, as we adjust $\xi_{k}$s flexibly based on the assigned weights. For the same reason, selecting the initial method—whether using the profile likelihood in the MLE or bootstrap samples in the LME—to select $\xi_{k}$ is also not crucial. A pilot simulation study in the Supplementary Material indicates that variations in the starting method (‘mle’ or ‘lme’) do not produce substantially different results. Similarly, selecting different confidence levels (90% or 95%) does not lead to substantial performance disparities. Therefore, in this study, we use the 95% confidence interval based on the profile likelihood (‘mle’ starter) of $\xi$ to select $\xi_{k}$.
Additionally, we considered a ‘conventional’ model averaging approach in which the parameters of each $GEV_{k}$ submodel are estimated using the MLE, and submodels are weighted based on smooth AIC scores. Here, the smooth AIC score (Claeskens & Hjort 2008) is defined as $\Delta AIC_{k}=AIC_{k}-min_{K}AIC_{k}$, so that the weight is obtained by:
$\hat{w}_{k}\ =\ {\frac{\exp(-\Delta AIC_{k}/2)}{\sum_{k=1}^{K}\exp(-\Delta AIC_{k}/2)}}.$ (24)
@@
For the LME, we utilized the ‘Lmoments’ or ‘lmomco’ packages in R, as provided by Karvanen (2006), based on the methods of Hosking (1990) and Asquith (2014). For the MLE, return level plots, and confidence interval based on profile likelihood, we primarily used the ‘ismev’ package (Coles 2001) in R. Additionally, for constrained optimization under (25), we employed the ‘Rsolnp’ package.
As presented in the Supplementary Material, a pilot simulation study revealed that MA methods exhibited substantial underestimation of return levels when $\xi\leq-0.3$. We applied a trimming approach to address this issue, removing the smallest observations. The abbreviations of the MA and other methods used in this study are listed in Table 1, where ‘trim=1’ and ‘trim=2’ indicate the cases where the minimum and second minimum observations, respectively, were excluded during estimation or weighting. Additionally, based on another pilot simulation study, we found that the weighting schemes ‘gLd’ and ‘med’ did not result in substantial performance differences. Therefore, for simplicity—given that 12 methods are already considered—the MA method with ‘med’ weighting is omitted from this simulation study.
Table 1: Abbreviations of the model averaging and other methods used in this study. Here, ‘gLd’ represents the generalized L-moments distance defined in (5), the weighting formula using smooth AIC is provided in (24), the constraints for Re.MLE is defined in (25), and the penalty function for MLE.CD is given in (27).

### v2

@@ 2.3 Estimation of high quantiles and building a surrogate model
which is a weighted average of the return level predictions from individual submodels. We aim to achieve this with lower estimation uncertainty than the MLE, LME, and other established methods.
The limitation of this MA approach is that it may not be useful for system description and understanding (Fletcher 2018). We consider constructing a weighted average (or ensemble) model from the $K$ GEV candidate models to address this limitation. However, the ensemble model may not necessarily follow a GEVD, even though each model $M_{k}$ does. Thus, simply taking a weighted sum of $K$ GEV candidate models is not an ideal solution. Instead, we propose building a surrogate model that still follows a GEVD and that best approximates the return level estimates generated by the MA method.
@@
where $\{r_{q_{i}}\}_{i=1,\dots,I}$ are the return levels corresponding to probability $q_{i}$, computed from a $GEVD(\mu,\sigma,\xi)$, and $\hat{r}_{q_{i}}$ are the corresponding return level estimates obtained using the MA method. To ensure a broad coverage of important quantiles, we selected $I=16$ probabilities ranging from $0.5$ to $0.999$. Although the surrogate model is only an approximation and may not be theoretically exact, it is computationally convenient and useful for further data analysis. This convenience includes additional predictions, model comparison, and model diagnostics such as quantile-per-quantile and return level plots. Of course, other models could also be used as surrogates instead of using the GEVD.
@@ 3.1 Conditional expectation
[…]
@@
[…]
@@
[…]
Since estimating $\xi$ is more challenging than estimating the other parameters of the GEVD and because $\hat{\xi}$ has the most substantial influence on the estimation of $r(\mu,\sigma,\xi)$—particularly when $\xi<-0.2$—conditioning on $\xi$ as a fixed value near $\hat{\xi}$ can substantially simplify the estimation process. The generally lower variability of a narrow submodel compared to a broader model is well demonstrated in Claeskens & Hjort (2008) under the local misspecification framework. Additionally, fitting submodels is computationally more efficient than fitting a full model.
[…]
@@
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\begin{pmatrix}\hat{\mu}_{2}(\xi),\hat{\sigma}_{2}(\xi)\text{ are two-dimentional }\\ \text{estimators for a given constant}\ \xi\end{pmatrix}$
$\ =\>\int r\{\hat{\mu}_{2}(\xi),\hat{\sigma}_{2}(\xi),\xi\}\times p\left[\,r\{\hat{\mu}_{2}(\xi),\hat{\sigma}_{2}(\xi),\xi\}\>|\>\hat{\xi}=\xi\,\right]\;dr$
@@
 $\ \xleftarrow{\text{ \large estimate }}\>\sum_{k=1}^{K}\>r\{\hat{\mu}_{2}(\xi_{k}),\hat{\sigma}_{2}(\xi_{k}),\xi_{k}\}\times\hat{w}(\xi_{k}),$
where $\xi_{k}$ values are selected based on the sampling distribution of $\xi$. Thus, our MA method can be interpreted as approximately integrating out the $\xi$ parameter.
@@ 3.2 Asymptotic distribution of MA return level
When submodel $M_{k}$ is fitted using the MLE, the MLE of $(\mu_{k},\ \sigma_{k})$ for a given $\xi_{k}$ follows an asymptotically normal distribution with mean $(\mu_{k},\;\sigma_{k})$ and covariance matrix $\Sigma_{k}$, which is the inverse of the Fisher information matrix, provided that $1<\xi_{k}<1/2$ (Bücher et al. 2021, for example). Consequently, $\hat{r}_{k}$ asymptotically follows a normal distribution with mean $r_{k}$ and variance:
@@
by the delta method, where $\nabla r_{k}$ is a gradient vector of $r_{k}$ with respect to $(\mu_{k},\;\sigma_{k})$. For the GEV distribution, the gradient vector of $r_{k}$ with respect to $(\mu_{k},\;\sigma_{k})$ is given by
$\nabla r_{k}\,=\,[1,\;\xi_{k}^{-1}(1-y_{p}^{\xi_{k}})]^{T},$ (14)
for given $\xi_{k}$, where $y_{p}=-\ln\>(1-p)$.
@@
Let $\hat{w}_{n}(k)$ and $\hat{r}_{n}(k)$ represent the weight and return level estimates obtained from data with sample size $n$. Additionally, let $v_{k}$ denote the asymptotic variance of $\hat{r}_{k}$, as provided in (13). Define $W_{k}$ as a random variable such that $\hat{w}_{n}(k)\xrightarrow{\text{ D ~}}W_{k}$ as $n\rightarrow\infty$, where $E(W_{k})=w_{k}$. Then we have the following result for the asymptotic distribution of the MA return level estimator:
@@
$\sum_{k=1}^{K}\hat{w}_{n}(k)\>\hat{r}_{n}(k)-r_{true}\>\xrightarrow{\text{ D ~}}\>\Lambda_{MA}\;=\;\sum_{k=1}^{K}w_{k}\>\Lambda_{k}$ (16)
where $\Lambda_{k}\,\sim N(0,\>v_{k})$.
Proof: Let $R_{k}$ be a normal random variable such that $\hat{r}_{n}(k)\xrightarrow{\text{ D ~}}R_{k}\sim N(r_{k,true},\>v_{k})$. Then, we have:
@@
Since $\hat{w}_{n}(k)$ is uniformly integrable, it follows that $E[\hat{w}_{n}(k)]\rightarrow w_{k}$, which implies that $\hat{w}_{n}(k)\xrightarrow{\text{ p ~}}w_{k}$ (Serfling 1980). Thus, the above equation simplifies to:
 $\ \xrightarrow{\text{ D ~}}\ \sum_{k=1}^{K}(w_{k}-W_{k})\;R_{k}\;+\;\sum_{k=1}^{K}W_{k}R_{k}\;-\;\sum_{k=1}^{K}w_{k}\;r_{k,true}$ (18)
@@
The Slutsky theorem and the continuity theorem of weak convergence are applied in the above derivation. This establishes the statement of the theorem.∎
@@
where $\underline{w}^{T}=(w_{1},\dots,w_{K})$ and $C$ is $Var(\underline{\Lambda})$ which is the asymptotic covariance matrix of $\hat{\>}\underline{r}$, where $\underline{\Lambda}^{T}=(\Lambda_{1},\dots,\Lambda_{K})$.
@@ 3.3 Asymptotic variance of MA return level
The diagonal elements of $C$, the asymptotic variances of $\hat{\>}\underline{r}$, are obtained using (13). However, the off-diagonal elements of $C$, representing the covariances $\text{Cov}(\hat{r}_{i},\;\hat{r}_{j})$ for $i\neq j$, are difficult to compute explicitly. To address this, we considered a variance formula directly from (8):
@@ 3.4 MA variance with random weights
For computing $Var(\Lambda_{MA})$ in real applications with data of size $n$, we may need to replace $\underline{w}$ in (19) with its estimate $\hat{\>}\underline{w}^{T}=(\hat{w}_{1},\dots,\hat{w}_{K})$. In this case, $\hat{\underline{w}}$ is a random vector derived from the sample, meaning that the formula (19) may need to be adjusted to account for the uncertainty in $\hat{\underline{w}}$. To calculate $Var(\Lambda_{MA})$ using the random vector $\hat{\underline{w}}$, we assume that $\hat{\underline{w}}$ and $\hat{\>}\underline{r}$ are independent. The expression for the variance is then given by:
@@
[…]
@@
[…]
We need to estimate E(w¯^)E($\hat$$\underline${w}}), $E(\hat{\>}\underline{r})$, and $D$ from the data for practical purposes. However, deriving these quantities theoretically from (7) is challenging. One possible approach is to use bootstrapping to approximate these quantities, although we did not adopt this method in our study. Instead, we assume a distribution for $\hat{\underline{w}}$. Given that $0\leq\hat{w}_{k}\leq 1$ for $k=1,\dots,K$ and that the weights satisfy $\sum_{k=1}^{K}\hat{w}_{k}=1$, it is reasonable to assume that $\hat{\underline{w}}$ follows a Dirichlet distribution with parameter vector $\underline{w}$, where $0\leq w_{k}\leq 1$ for $k=1,\dots,K$ and $\sum_{k=1}^{K}w_{k}=1$. Then, under this assumption,
@@
for all $i\neq j$. These quantities form the elements of $D$. The covariance matrix $D$ can then be approximated by substituting $\hat{\>}\underline{w}$ in place of $\underline{w}$, yielding the estimator $\hat{D}$.
To approximate $E(\hat{\>}\underline{r})$, we used a moving average of order $q$ for $\hat{\>}\underline{r}$ instead of directly substituting $\hat{\>}\underline{r}$ into $E(\hat{\>}\underline{r})$. With these approximations, we obtain $Var(\hat{r}_{\text{MA}})$ in (22) using the following formulation:
@@
where $\widetilde{\hat{\>}\underline{r}(q)}^{T}$ is the order $q$ moving average of $\hat{\>}\underline{r}^{T}=(\hat{r}_{1},\dots,\hat{r}_{K})$. We chose $q=3$ in this study. This formula is computationally much faster than using the bootstrap to estimate E(w¯^)E($\hat$$\underline${w}}), $E(\hat{\>}\underline{r})$, and $D$.
@@ 4.1 Computing strategy
Selecting the number ($K$) of candidate models is an important step in the proposed method. One recommended strategy is to start with a relatively large $K$ (e.g., $K=18$) and then remove submodels with small weights, such as those with $w_{k}\leq 0.01$. Additionally, if the weight assigned to either the smallest $\xi_{k}$ or the largest $\xi_{k}$ remains above a certain threshold (e.g., 0.1), the method adds a few additional $\xi_{k}$ values beyond the lower or upper bounds. The weights are then recalculated using the updated number of submodels, denoted by $K^{\prime}$, and the MA return level estimate $\hat{r}_{MA}$ is obtained using the new weight vector ${\>\hat{\underline{w}}^{\prime}}^{T}=(\hat{w}^{\prime}_{i_{1}},\dots,\hat{w}^{\prime}_{i_{K^{\prime}}})$. Thus, choosing the number $K$ may not be so critical provided that $K$ is not too small. We would recommend $K$ to be greater than or equal to seven, according to our experience. Nevertheless, in Section 6, we consider a systematic way to select an appropraite $K$.
Thus, selecting between 90% and 95% confidence intervals for choosing $\xi_{k}$ values may no longer be critically important, as we adjust $\xi_{k}$s flexibly based on the assigned weights. For the same reason, selecting the initial method—whether using the profile likelihood in the MLE or bootstrap samples in the LME—to select $\xi_{k}$ is also not crucial. A pilot simulation study in the Supplementary Material indicates that variations in the starting method (‘mle’ or ‘lme’) do not produce substantially different results. Similarly, selecting different confidence levels (90% or 95%) does not lead to substantial performance disparities. Therefore, in this study, we use the 95% confidence interval based on the profile likelihood (‘mle’ starter) of $\xi$ to select $\xi_{k}$.
Additionally, we considered a ‘conventional’ model averaging approach in which the parameters of each $GEV_{k}$ submodel are estimated using the MLE, and submodels are weighted based on smooth AIC scores. Here, the smooth AIC score (Claeskens & Hjort 2008) is defined as $\Delta AIC_{k}=AIC_{k}-min_{1\leq k\leq K}AIC_{k}$, so that the weight is obtained by:
$\hat{w}_{k}\ =\ {\frac{\exp(-\Delta AIC_{k}/2)}{\sum_{j=1}^{K}\exp(-\Delta AIC_{j}/2)}}.$ (25)
@@
For the LME, we utilized the ‘Lmoments’ or ‘lmomco’ packages in R, as provided by Karvanen (2006), based on the methods of Hosking (1990) and Asquith (2014). For the MLE, return level plots, and confidence intervals based on profile likelihood, we primarily used the ‘ismev’ package (Coles 2001) in R. Additionally, for constrained optimization under (26), we employed the ‘Rsolnp’ package.
As presented in the Supplementary Material, a pilot simulation study revealed that MA methods exhibited substantial underestimation of return levels when $\xi\leq-0.3$. We applied a trimming approach to address this issue by removing the smallest observations from the sample. The abbreviations of the MA and other methods used in this study are listed in Table 1, where ‘trim=1’ and ‘trim=2’ indicate the cases where the minimum and second minimum observations, respectively, were excluded during estimation or weighting. Additionally, based on another pilot simulation study, we found that the weighting schemes ‘gLd’ and ‘med’ did not result in substantial performance differences. Therefore, for simplicity—given that a total of 12 methods are already considered—the MA method with ‘med’ weighting is omitted from this simulation study.
Table 1: Abbreviations of the model averaging and other methods used in this study. Here, ‘gLd’ represents the generalized L-moments distance defined in (5), the weighting formula using smooth AIC is provided in (25), the constraints for Re.MLE are defined in (26), and the penalty function for MLE.CD is given in (28).